a commitee of three teachers and three students is being formed. There are 100 students (80 of which are women) and 20 teachers (ten are women). What are the chances that the committee will be made up of all women?
I said there are 90 women all together and 30 men therefore there is a 60% chance that the committee will be all women

Answer by stanbon(75887)

a commitee of three teachers and three students is being formed.
There are 100 students (80 of which are women) and 20 teachers (ten are women). What are the chances that the committee will be made up of all women?
-------------------------------
P(3 students are women) = 80C3/100C3 = 0.5081
---
P(3 teachers are women) = 10C3/20C3 = 0.1053
--------------------------------------------------
P(committee of 6 women) = 0.5081*0.1053 = 0.0535
=========================================================
Cheers,
Stan H.